The productive capability of an economy is such that to produce 5 units of military good it takes 2 workers to be employed while 10 units of consumer goods require 3 workers. Resources are limited in such a way that only 75 units of military good can be produced when all resources are employed.

a) How much workers are in the economy?

b) What is the maximum amount of consumer goods that can be produced?

c) Construct the production possibility schedule and curve for this economy.


d) Graphically represent what would happen if 12 additional workers were added

to the economy.

e) Graphically represent what would happen if the productivity of the workers

were reduced to 2 units of military good and 5 units of consumer goods.

f) Graphically represent what would happen if productivity for military goods

remained the same but it now required 2 workers to produce 10 units of consumer good


1
Expert's answer
2021-05-18T12:15:39-0400

a)

The number of military goods per worker that could be produced is


"\frac{number\ of\ millitary\ goods\ produced}{numbers\ of\ workers\ taken\ to\ produce\ it}"


"\frac{5units}{2units}"


Each worker receives 2.5 units of military goods.


The number of people engaged in the production process will be,


"\frac{Number\ of\ military\ goods\ produced}{Number\ of\ millitary\ goods\ produced\ per\ worker}"


"\frac{75units}{2.5units}"

​

30 Workers.


The number of consumer goods produced per worker.


"\frac{Number \ of\ consumer\ goods\ produced}{Number\ of\ workers\ taken\ to \ produce\ it}"


"\frac{10units}{3workers}"


The units of consumer goods per worker will be 3.33.


b) The number of consumer goods that will be produced.


Numbers of workers used for production "\times" Number of consumer goods produced per worker.


"=30\times3.33"


"=99.99"


By rounding this off we get,


"100\ units\ of\ consumer\ goods\ produced."


c)

The different combinations of items that can be created in the economy utilizing the resources and technology available would be the consequence of the production possibility frontier. There are a total of 30 people in the economy, resulting in the production of 75 units of military goods or 100 units of consumer goods.




d) The addition of 12 workers to the labor force would bring the overall number of workers to 42, increasing the production of both military and consumer items. As a result, the ppf would shift from AB to AB'.




e) As productivity declines, the quantity of military items per worker drops to two per worker, but consumer goods rise to five units per worker, resulting in an unequal shift in the production possibility frontier. As a result, the ppf shifts from AB to AB" generating 60 military items and 150 consumer items.

f)

As productivity changes,

2 workers = 10 units of consumer goods, 

5 units per worker

150 units = 30 workers

As a result, on the y-axis, where we have consumer goods, the production possibility frontier will shift outward.